# Backdoor Attacks against No-Reference Image Quality Assessment Models via A Scalable Trigger

### v1

@@ Introduction
Alongside the impressive performance of DNNs, concerns about their security have grown (Yu et al. 2022, 2024b; Xia et al. 2024a, b; Wang et al. 2023, 2024a, 2024b, 2024c; Lin et al. 2024; Meng et al. 2024). Adversarial attacks (AA) on NR-IQA models have recently received considerable attention. These attacks aim to substantially alter predicted scores with imperceptible perturbations applied to input images. Despite the insights provided by these attacks in NR-IQA models, they exhibit several intrinsic shortcomings: 1) Some attacks (Zhang et al. 2022) assume a white-box scenario, wherein the attacker has full access to the model and its parameters, limiting their practical utility in the real world. 2) Certain attacks (Korhonen and You 2022), relying on surrogate models to generate adversarial examples and transfer them to the target model, often suffer from reduced effectiveness in limited transferability. 3) The generation of adversarial examples is formulated as an optimization problem, requiring substantial computational resources and time to iteratively optimize solutions for each sample. 4) Most attacks (Zhang et al. 2022, 2023) aim to create untargeted attacks, focusing on inducing significant deviations rather than shifting predictions to a specific target.
@@
To tackle these challenges, we consider an alternative threat: backdoor attacks (BA), which are more efficient and require no white-box access to the model during testing. In this scenario, the adversary can inject a stealthy backdoor into the target model corresponding to a unique trigger during training (Fang and Choromanska 2022; Yu et al. 2023, 2024a, 2024c; Liu et al. 2024b; Zheng et al. 2024), enabling the compromised model to function normally on benign samples but exhibit malicious behavior when presented with samples containing the trigger during inference. Data poisoning is the prevailing method for executing BA (Chen et al. 2017), whereby the adversary compromises specific training samples to embed the backdoor. Models trained on poisoned data with the trigger, consistently predicted a specific target whenever the trigger is present during testing. These attacks are challenging to detect, as the backdoored models maintain high performance on clean test data. Current poisoning-based BA falls into two categories: 1) poison-label attacks, which contaminate training examples and alter their labels to the target value; and 2) clean-label attacks, which solely affect training examples while maintaining their labels.
@@
Image Quality Assessment. IQA tasks aim at predicting image quality scores that are consistent with human perception, often represented by Mean Opinion Score (MOS). These tasks can be categorized into Full-Reference (FR) and No-Reference (NR). In FR-IQA, the objective is to compare a distorted image with its reference image to predict the quality score of the distorted image. However, obtaining reference images can be challenging in some real-world scenarios, leading to the emergence of NR-IQA, which aims to predict the quality score using only the distorted image. Some approaches (Mittal, Moorthy, and Bovik 2012; Ghadiyaram and Bovik 2017) leverage hand-crafted features derived from Natural Scene Statistics, while others explore the influence of semantic information on perceived image quality. For instance, Hyper-IQA (Su et al. 2020) employs a hypernetwork to obtain different quality estimators for images with varying content. DBCNN (Zhang et al. 2020) utilizes two independent neural networks to extract distorted and semantic information from images, which are then combined using bilinear pooling. Additionally, various studies have investigated the effectiveness of architectures in NR-IQA. TReS (Golestaneh, Dadsetan, and Kitani 2022) and MUSIQ (Ke et al. 2021) leverage vision transformers (Dosovitskiy et al. 2021) and demonstrate their efficacy in NR-IQA tasks. While adversarial attacks against IQA tasks have received some attention (Zhang et al. 2022; Shumitskaya, Antsiferova, and Vatolin 2022; Liu et al. 2024a; Zhang et al. 2023), a more concerning and practical threat, known as backdoor attacks, remains unexplored.
@@
Backdoor Attacks. Both the BA (Chen et al. 2017) and AA (Szegedy et al. 2014) intend to modify the benign samples to mislead the DNNs, but they have some intrinsic differences. At the inference stage, AA (Madry et al. 2018; Ilyas et al. 2018) require much computational resources and time to generate the perturbation through iterative optimizations, and thus are not efficient in deployment. However, the perturbation (trigger) is known or easy to generate for BA. From the perspective of the attacker’s capacity, BA have access to poisoning training data, which adds an attacker-specified trigger (e.g. a local patch) and alters the corresponding label. BA on DNNs have been explored in BadNet (Gu, Dolan-Gavitt, and Garg 2017) for image classification by poisoning some training samples, and the essential characteristic consists of 1) backdoor stealthiness, 2) attack effectiveness on poisoned images, 3) low performance impact on clean images. Based on the capacity of attackers, BA can be categorized into poisoning-based and non-poisoning-based attacks (Li et al. 2020b). For poisoning-based attacks (Li et al. 2020a, 2021), attackers can only manipulate the dataset by inserting poisoned data, and have no access to the model training process. In contrast, non-poisoning-based attacks (Dumford and Scheirer 2020; Rakin, He, and Fan 2020; Guo, Wu, and Weinberger 2020; Doan et al. 2021) inject the backdoor by modifying the model parameters instead. For trigger generations, most attacks (Chen et al. 2017; Steinhardt, Koh, and Liang 2017) rely on fixed triggers, and several recent methods (Nguyen and Tran 2020; Liu et al. 2020) extend it to be sample-specific. For the trigger domain, several works (Zeng et al. 2021; Wang et al. 2022; Yue et al. 2022) consider the trigger in the frequency domain due to its advantages (Cheng et al. 2023). FTrojan (Wang et al. 2022) blockifies images and adds the trigger in the DCT domain, but it selects two fixed channels with fixed magnitudes.
@@
Our Backdoor’s Goals. Unlike traditional BA on classifiers, where the output space is discrete, the output space of IQA is continuous. Hence, our goal is to embed a backdoor capable of manipulating the model to output any values, rather than only one target class typically seen in BA on classifiers. This ensures that any input $\bm{x}\!\in\!\mathcal{X}$ with MOS value $y$, once altered with the trigger $\bm{t}$ into the poisoned data $T(\bm{x},{\alpha}\!\cdot\!\bm{t})$, is incorrectly predicted into a specific target value $y\!+\!\alpha\!\cdot\!{\Delta{y}}_{t}$. Here, $\alpha\!\!\in\!\![-1,1]$ controls the variations, and regulates the magnitude of the trigger $\bm{t}$. $T$ is the trigger injection function, and $\alpha\Delta{y}_{t}$ is the target variation. We can find some theoretical support, since when using the Taylor expansion and considering the first-order derivative, the output of $f_{\bm{\theta}}$ can exhibit a certain degree of linearity with respect to $\alpha$, as follows:
$\begin{split}f_{\bm{\theta}}\!(T(\bm{x},{\alpha}\!\cdot\!\bm{t}))\!=\!f_{\bm{\theta}}\!(T(\bm{x},0\!\cdot\!\bm{t}))\!+\!\alpha\frac{\partial f_{\bm{\theta}}\!(T(\bm{x},\alpha_{0}\!\cdot\!\bm{t}))}{\partial\alpha_{0}}\Big{|}_{\alpha_{0}\!=\!0}.\end{split}$ (1)
@@
Input : Subset $\mathcal{D}_{s}=\{(\bm{x}_{i},{y}_{i})\}_{i=1}^{N_{s}}$, surrogate model $f_{\bm{\theta}_{s}}$, perturbation bound $\epsilon$, epochs $e_{1},e_{2}$
Output : Optimized trigger $\bm{t}$
$\bm{\theta}_{s}\leftarrow\text{train}~{}f_{\bm{\theta}_{s}}~{}\text{on}~{}\mathcal{D}_{s}~{}\text{by}~{}\mathcal{L}_{1}~{}\text{and Adam for }e_{1}~{}\text{epochs}$;
for $i\leftarrow 1$ to $e_{2}$  do
       for $\{\bm{x},{y}\}\in\mathcal{D}_{s}$ do
             $\mathcal{L}_{UAP}=-\underbrace{\lVert f_{\bm{\theta}_{s}}(T\left(\bm{x},\bm{t}\right))-f_{\bm{\theta}_{s}}(\bm{x})\lVert_{1}}_{\text{effectiveness of UAP-DCT}}+\lambda\cdot\underbrace{\text{max}(\mathcal{L}_{mse}(\bm{x},T(\bm{x},\bm{t})),\epsilon^{2})}_{\text{invisibility of UAP-DCT}}$;
             $\bm{t}\leftarrow\text{minimize}~{}\mathcal{L}_{UAP}~{}\text{using Adam optimizer}$;
       end for
end for
Algorithm 1 Searching Universal Adversarial Perturbation in the 2d-DCT domain (UAP-DCT)
Exploiting DCT Space Vulnerability. Indeed, research has revealed that state-of-the-art deep neural network classifiers are vulnerable to a phenomenon known as Universal Adversarial Perturbations (UAP), which refers to a universal and invisible perturbation vector that can cause high-probability misclassifications of natural images. Inspired by the efficacy of universally applicable perturbations in misleading well-trained models, we leverage UAP in the DCT space, denoted as UAP-DCT, as the trigger $\bm{t}$. By incorporating these UAP-DCT triggers into a poisoned dataset, the attacker can amplify the susceptibility of a trained IQA model to such triggers, thereby enhancing attack performance and facilitating more effective manipulation of the attacked outputs. Algorithm 1 outlines the optimization process for the trigger $\bm{t}$, involving two loss terms. The first loss aims to generate effective UAP-DCT, while the second loss regulates the magnitude of UAP-DCT to ensure the induced perturbations remain invisible in the spatial domain, thereby enhancing the imperceptibility of the poisoned images.
@@
During the inference, for any test sample $(\bm{x},y)\sim\mathcal{C}$, the adversary can trigger the backdoor using the poisoned image $T\left(\bm{x},{\alpha}\cdot\bm{t}\right)$. The adversary is free to choose the scaling coefficient $\alpha$, thus manipulating the output to the desired target $f_{\bm{\theta}^{*}}(\bm{x})+\alpha\cdot{\Delta{y}}_{t}$ according to their intentions.
@@
$\begin{split}&\mathbb{P}(\tilde{y}|\bm{x},\!\alpha)\!=\!\mathcal{N}(\!\tilde{y};y\!+\!\alpha{\Delta{y}}_{t},\!\sigma^{2}),\mathbb{P}(\tilde{y}|\bm{x})\!=\!\mathcal{N}(\!\tilde{y};y,\!\sigma^{2}),\\ &\mathbb{P}(\tilde{y})=\mathcal{N}(\tilde{y};\mu_{\tilde{y}},\sigma^{2}),\mathbb{P}(\tilde{y}|\alpha)=\mathcal{N}(\tilde{y};\mu_{\tilde{y}|\alpha},\sigma^{2}),\end{split}$ (3)
@@
$\begin{split}\!\!\mu_{\tilde{y}}\!=\!\mathbb{E}[\tilde{y}]\!=\!\!\!\!\mathop{\mathbb{missing}}{E}\limits_{y\sim\mathcal{D}_{t}}\![y]\!=\!\mu_{y},\mu_{\tilde{y}|\alpha}\!=\!\!\!\!\mathop{\mathbb{missing}}{E}\limits_{y\sim\mathcal{D}_{t}}\![y]\!+\!\alpha{\Delta{y}}_{t}\!=\!\mu_{y}\!+\!\alpha{\Delta{y}}_{t}.\end{split}$ (5)
@@
Remark 1 outlines the sampling strategy for $\alpha$ based on Assumption 1. Additionally, for Eq. 3, it presents the mathematical relationship $\mathbb{P}(\tilde{y}|\bm{x},\alpha)\mathbb{P}(\tilde{y})=\mathbb{P}(\tilde{y}|\bm{x})\mathbb{P}(\tilde{y}|\alpha)$ for any test data. Given that the model $f_{\bm{\theta}}$ learns this property from the training set, it is crucial that the training set adheres to this equation. To increase the likelihood of this equation holding in the training set and ensure consistency between training on $\mathcal{D}_{p}$ and evaluation during the attack phase, we must consider whether modifying $\bm{x}$ to $\bm{x^{\prime}}$ is necessary. This leads to the development of the following theorem.
Input : Benign dataset $\mathcal{D}=\{(\bm{x}_{i},{y}_{i})\}_{i=1}^{N}$, surrogate model $f_{\bm{\theta}_{s}}$, poisoning ratio $r$, hyperparameter for UAP-DCT ($\epsilon$, $e_{1},e_{2}$), hyperparameter for TAEs (target $\mu_{y}$,bound $\epsilon_{t}$, step size $\alpha_{t}$, iterations $I_{t}$).
Output : Optimized trigger $\bm{t}$, Modified dataset $\mathcal{D}_{t}$.
$\mathcal{D}_{s}\leftarrow~{}\text{randomly sample a subset from}~{}\mathcal{D}~{}\text{with ratio}~{}r$;
# Train surrogate model and search trigger
[…]
for $\{\bm{x}_{i},{y_{i}}\}\in\mathcal{D}_{s}$ do
       $\alpha_{i}=(y_{i}-\mu_{y})/{{\Delta{y}}_{t}}$;
       $\bm{x^{\prime}}_{i}\leftarrow$ targeted-PGD(model=$f_{\bm{\theta}_{s}}$, input=$\bm{x}_{i}$, target=$\mu_{y}$, bound=$\epsilon_{t}$, step=$\alpha_{t}$, iter=$I_{t}$);
end for
Return $\mathcal{D}_{p}=\{(T(\bm{x^{\prime}}_{i},g(\alpha_{i})\cdot\bm{t}),y_{i}+\alpha_{i}{\Delta{y}}_{t})\}_{i=1}^{N_{s}}$, $\mathcal{D}_{c}=\mathcal{D}\backslash\mathcal{D}_{s}$, $\mathcal{D}_{t}=\mathcal{D}_{c}\cup\mathcal{D}_{p}$.
@@
Table 2: Imperceptibility of the poisoned data in terms of $\text{PSNR}_{1}\uparrow$ for all competing methods.
@@
Datasets and Models. In this paper, we choose the widely used LIVEC dataset (Ghadiyaram and Bovik 2015) consisting of 1162 images with sizes $500\times 500$ and KonIQ-10k dataset (Hosu et al. 2020) consisting of 10073 images with sizes $512\times 384$. For LIVEC, consistent with prior research (Liu et al. 2024a), we randomly select 80% of the images as the training set, and the remaining 20% as the testing set. For KonIQ-10k, we adopt the official split with 70% of the images as the training set. We assess the effectiveness of our attacks against two CNN-based NR-IQA models: HyperIQA (Su et al. 2020), and DBCNN (Zhang et al. 2020), and one Transformer-based TReS (Golestaneh, Dadsetan, and Kitani 2022). The training configurations are in the appendix.
Baseline Attack Methods. We compare our attacks with existing poisoning-based BA. We utilize Blended (Chen et al. 2017), WaNet (Nguyen and Tran 2021), and FTrojan (Wang et al. 2022), which are representative non-patch-based invisible attacks. For all baselines, we employ the coefficient $\alpha$ to adjust the intensity and strength of the triggers, following a similar approach to our proposed attacks. For FTrojan, we incorporate the triggers into the same mid-frequency channels. We apply the same configurations to all methods, with the sole exception of the trigger. More details are in the appendix.
Evaluation Metrics. For the evaluation on benign data, we follow previous works and consider three metrics, i.e., RMSE, PLCC, and SROCC, defined in Appendix. To evaluate the attack effectiveness, we include the mean absolute error (MAE), and the mean ratio of amplification (MRA) across the test set regarding each $\alpha\!\!\in\!\![-1,1]$. While it is impossible to evaluate the overall attack success by considering all values of $\alpha$, we estimate it using finite $\alpha$ from a selected set $A$. Therefore, we utilize the mean of those metrics for an overall evaluation:
$\begin{split}&\text{MAE}(\alpha)\!\!=\!\!\!\!\!\!\!\mathop{\mathbb{missing}}{E}\limits_{(\bm{x},y)\sim\mathcal{C}}\!\!\Big{[}\!\big{\lvert}\Delta_{{a}}(\bm{x},\alpha)\!\!-\!\!\Delta_{{t}}(\alpha)\big{\lvert}\!\Big{]}\!,\text{MRA}(\alpha)\!\!=\!\!\!\!\!\!\!\mathop{\mathbb{missing}}{E}\limits_{(\bm{x},y)\sim\mathcal{C}}\!\!\Big{[}\!\frac{\Delta_{{a}}(\bm{x},\alpha)}{\Delta_{{t}}(\alpha)}\!\Big{]}\!,\\ &\text{mMAE}\!=\!\frac{1}{n(A)}\!\sum_{\alpha\in A}\!\text{MAE}(\alpha),\text{mMRA}\!=\!\frac{1}{n(A)}\!\sum_{\alpha\in A}\!\text{MRA}(\alpha),\\ &\text{where}~{}\Delta_{{a}}(\bm{x},\alpha)\!=\!f_{\bm{\theta}^{*}}(T(\bm{x},{\alpha}\cdot\bm{t}))\!-\!f_{\bm{\theta}^{*}}(\bm{x}),\Delta_{{t}}(\alpha)\!=\!\alpha\cdot{\Delta{y}}_{t}.\end{split}$
@@
Settings. Since the MOS range for both datasets are $[0,100]$, we set the maximum deviation $\Delta{y_{t}}\!\!=\!\!40$, ensuring this deviation is sufficient to yield potent attacks. To regulate the imperceptibility of poisoned images, in Algorithm 1, we set $\lambda\!\!=\!\!10^{8}$, $e_{1}\!\!=\!\!24$, $e_{2}\!\!=\!\!50$, and the bound $\epsilon\!\!=\!\!8/255$ for LIVEC. However, for KonIQ-10k, where a larger number of images are eligible for poisoning, we adjust $\epsilon$ to a lower value $4/255$. In the main experiments, we set the poisoning ratio $r\!\!=\!\!20\%$ across all attacks and for both datasets, and opt for the HyperIQA as the surrogate model. For the sampling strategy of $\alpha$ to construct the poisoned dataset $\mathcal{D}_{p}$, we first consider sampling from {±$\{\pm 1,\pm\frac{3}{4},\pm\frac{1}{2},\pm\frac{1}{4}\}$ with probabilities $\{0.4,0.3,0.2,0.1\}$, and provide results of other strategies in the ablation study.
@@
In addition, we provide the $\text{MRA}(\alpha)$ in Fig. 2(a). We can see that our method is capable of achieving a significant deviation when $\alpha$ has an absolute value greater than 0.5. However, when $\alpha$ falls within [-0.5, 0.5], the manipulation becomes less precise, as the model finds it more challenging to recognize the trigger. However, we believe this limitation is tolerable given that attackers generally aim for significant shifts in outputs rather than minor alterations. Results of $\text{MAE}(\alpha)$, visualized results of poisoned images, and results of P-BAIQA with lower poisoning ratios are in appendix.

### v3

@@ Introduction
Alongside the impressive performance of DNNs, concerns about their security have grown (Liang et al. 2021; Yu et al. 2022, 2024b; Xia et al. 2024a, b; Wang et al. 2024a, b). Adversarial attacks (AA) on NR-IQA models have recently received considerable attention. These attacks aim to substantially alter predicted scores with imperceptible perturbations applied to input images. Despite the insights provided by these attacks in NR-IQA models, they exhibit several intrinsic shortcomings: 1) Some attacks (Zhang et al. 2022) assume a white-box scenario, wherein the attacker has full access to the model and its parameters, limiting their practical utility in the real world. 2) Certain attacks (Korhonen and You 2022), relying on surrogate models to generate adversarial examples and transfer them to the target model, often suffer from reduced effectiveness in limited transferability. 3) The generation of adversarial examples is formulated as an optimization problem, requiring substantial computational resources and time to iteratively optimize solutions for each sample. 4) Most attacks (Zhang et al. 2022, 2023) aim to create untargeted attacks, focusing on inducing significant deviations rather than shifting predictions to a specific target.
@@
To tackle these challenges, we consider an alternative threat: backdoor attacks (BA), which are more efficient and require no white-box access to the model . In this scenario, the adversary can inject a stealthy backdoor into the target model corresponding to a unique trigger during training (Saha, Subramanya, and Pirsiavash 2020; Fang and Choromanska 2022; Yu et al. 2023, 2024a, 2024c; Liu et al. 2024b; Zheng et al. 2024), enabling the compromised model to function normally on benign samples but exhibit malicious behavior when presented with samples containing the trigger during inference. Data poisoning is the prevailing method for executing BA (Chen et al. 2017), whereby the adversary compromises specific training samples to embed the backdoor. Models trained on poisoned data with the trigger, consistently predicted a specific target whenever the trigger is present during testing. These attacks are challenging to detect, as the backdoored models maintain high performance on clean test data. Current poisoning-based BA falls into two categories: 1) poison-label attacks, which contaminate training examples and alter their labels to the target value; and 2) clean-label attacks, which solely affect training examples while maintaining their labels.
@@
Image Quality Assessment. IQA tasks aim at predicting image quality scores that are consistent with human perception, often represented by Mean Opinion Score (MOS). These tasks can be categorized into Full-Reference (FR) and No-Reference (NR). In FR-IQA, the objective is to predict the quality score of the distorted image by comparing with its reference image. However, obtaining reference images can be challenging in real-world scenarios, leading to the emergence of NR-IQA, which predicts using only the distorted image. Some approaches (Mittal, Moorthy, and Bovik 2012; Ghadiyaram and Bovik 2017) leverage hand-crafted features, while others explore the influence of semantic information. For instance, Hyper-IQA (Su et al. 2020) employs a hypernetwork to obtain different quality estimators for images with varying content. DBCNN (Zhang et al. 2020) utilizes two independent neural networks to extract distorted and semantic information from images, which are then combined using bilinear pooling. Additionally, various studies have investigated the effectiveness of architectures in NR-IQA. TReS (Golestaneh, Dadsetan, and Kitani 2022) and MUSIQ (Ke et al. 2021) leverage vision transformers (Dosovitskiy et al. 2021) and demonstrate their efficacy in NR-IQA tasks. While adversarial attacks against IQA tasks have received some attention (Zhang et al. 2022; Shumitskaya, Antsiferova, and Vatolin 2022; Liu et al. 2024a; Zhang et al. 2023), a more concerning and practical threat, known as backdoor attacks, remains unexplored.
@@
Backdoor Attacks. BA (Chen et al. 2017) and AA (Szegedy et al. 2013) intend to modify the benign samples to mislead the DNNs, but they have some intrinsic differences. At the inference stage, AA (Madry et al. 2018; Ilyas et al. 2018) require much computational resources and time to generate the perturbation through iterative optimizations, and thus are not efficient in deployment. However, the perturbation (trigger) is known or easy to generate for BA. From the perspective of the attacker’s capacity, BA have access to poisoning training data, which adds an attacker-specified trigger (e.g. a local patch) and alters the corresponding label. BA on DNNs have been explored in BadNet (Gu, Dolan-Gavitt, and Garg 2017) for image classification by poisoning some training samples, and the essential characteristic consists of 1) backdoor stealthiness, 2) attack effectiveness on poisoned images, 3) low performance impact on clean images. Based on the capacity of attackers, BA can be categorized into poisoning-based and non-poisoning-based attacks (Li et al. 2020b). For poisoning-based attacks (Li et al. 2020a, 2021), attackers can only manipulate the dataset by inserting poisoned data, and have no access to the model training process. In contrast, non-poisoning-based attacks (Dumford and Scheirer 2020; Rakin, He, and Fan 2020; Guo, Wu, and Weinberger 2020; Doan et al. 2021) inject the backdoor by modifying the model parameters instead. For trigger generations, most attacks (Chen et al. 2017; Steinhardt, Koh, and Liang 2017) rely on fixed triggers, and several recent methods (Nguyen and Tran 2020; Liu et al. 2020) extend it to be sample-specific. For the trigger domain, several works (Zeng et al. 2021; Wang et al. 2022; Yue et al. 2022) consider the trigger in the frequency domain due to its advantages (Cheng et al. 2023). FTrojan (Wang et al. 2022) blockifies images and adds the trigger in the DCT domain, but it uses two fixed channels with fixed magnitudes.
@@
Our Backdoor’s Goals. Unlike traditional BA on classifiers, where the output space is discrete, the output space of IQA is continuous. Hence, our goal is to embed a backdoor capable of manipulating the model to output any values, rather than only one target class typically seen in BA on classifiers. This ensures that any input $\bm{x}\!\in\!\mathcal{X}$ with MOS value $y$, once altered with the trigger $\bm{t}$ into the poisoned data $T(\bm{x},{\alpha}\!\cdot\!\bm{t})$, is incorrectly predicted into a specific target value $y\!+\!\alpha\!\cdot\!{\Delta{y}}_{t}$. Here, $\alpha\!\!\in\!\![-1,1]$ controls the variations, and regulates the magnitude of $\bm{t}$. $T$ is the trigger injector, and $\alpha\cdot\Delta{y}_{t}$ is the target variation. We can see that when using the Taylor expansion and considering the first-order derivative, the output of $f_{\bm{\theta}}$ has a certain degree of linearity with respect to $\alpha$, as follows:
$\begin{split}f_{\bm{\theta}}(T(\bm{x},{\alpha}\cdot\bm{t}))=f_{\bm{\theta}}(T(\bm{x},0\cdot\bm{t}))+\alpha\frac{\partial f_{\bm{\theta}}(T(\bm{x},\alpha_{0}\cdot\bm{t}))}{\partial\alpha_{0}}\Big{|}_{\alpha_{0}=0}.\end{split}$
@@
Algorithm 1 Searching UAP in the DCT domain (UAP-DCT)
0:  Subset $\mathcal{D}_{s}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N_{s}}$, surrogate model $f_{\bm{\theta}_{s}}$, perturbation bound $\epsilon$, epochs $e_{1}$, $e_{2}$
0:  Optimized trigger $\bm{t}$
1:  Train $f_{\bm{\theta}_{s}}$ on $\mathcal{D}_{s}$ using loss $\mathcal{L}_{1}$ and Adam for $e_{1}$ epochs
2:  for $i=1$ to $e_{2}$ do
3:     for each $(\bm{x},y)\in\mathcal{D}_{s}$ do
4:        $\mathcal{L}_{UAP}=-\underbrace{\lVert f_{\bm{\theta}_{s}}(T(\bm{x},\bm{t}))-f_{\bm{\theta}_{s}}(\bm{x})\rVert_{1}}_{\text{effectiveness of UAP-DCT}}$   $\quad\quad\quad+\lambda\cdot\underbrace{\max(\mathcal{L}_{mse}(\bm{x},T(\bm{x},\bm{t})),\epsilon^{2})}_{\text{invisibility of UAP-DCT}}$
5:        Update $\bm{t}$ by minimizing $\mathcal{L}_{UAP}$ using the Adam
6:     end for
7:  end for
Exploiting DCT Space Vulnerability. Indeed, research has revealed that state-of-the-art DNN classifiers are vulnerable to a phenomenon known as Universal Adversarial Perturbations (UAP), which refers to a universal and invisible perturbations that can cause high-probability misclassifications. Inspired by the efficacy of universally applicable perturbations in misleading well-trained models, we leverage UAP in the DCT space, denoted as UAP-DCT, as the trigger $\bm{t}$. By incorporating these UAP-DCT triggers into a poisoned dataset, the attacker can amplify the susceptibility of a trained IQA model to such triggers, thereby enhancing attack performance and facilitating more effective manipulation of the attacked outputs. Algorithm 1 outlines the optimization process for the trigger $\bm{t}$, involving two loss terms. The first loss aims to generate effective UAP-DCT, while the second loss regulates the magnitude of UAP-DCT to ensure the induced perturbations remain invisible in the spatial domain, thereby enhancing the imperceptibility of the poisoned images.
@@
During the inference, for any test sample $(\bm{x},y)$, the adversary can trigger the backdoor using the poisoned $T\left(\bm{x},{\alpha}\cdot\bm{t}\right)$. The adversary is free to choose the $\alpha$, thus manipulating the output to the desired target $f_{\bm{\theta}^{*}}(\bm{x})+\alpha\cdot{\Delta{y}}_{t}$ as intended.
@@
$\begin{split}&\mathbb{P}(\tilde{y}|\bm{x},\alpha)=\mathcal{N}(\tilde{y};y+\alpha{\Delta{y}}_{t},\sigma^{2}),~{}\mathbb{P}(\tilde{y}|\bm{x})=\mathcal{N}(\tilde{y};y,\sigma^{2}),\\ &\mathbb{P}(\tilde{y})=\mathcal{N}(\tilde{y};\mu_{\tilde{y}},\sigma^{2}),~{}\mathbb{P}(\tilde{y}|\alpha)=\mathcal{N}(\tilde{y};\mu_{\tilde{y}|\alpha},\sigma^{2}),\end{split}$ (2)
@@
$\begin{split}\mu_{\tilde{y}}=\mathop{\mathbb{missing}}{E}\limits_{y\sim\mathcal{D}_{t}}[y]=\mu_{y},~{}\mu_{\tilde{y}|\alpha}=\mu_{\tilde{y}}+\alpha{\Delta{y}}_{t}=\mu_{y}+\alpha{\Delta{y}}_{t}.\end{split}$ (4)
@@
Remark 1 outlines the sampling strategy for $\alpha$ based on Assumption 1. Additionally, for Eq. 2, it presents the mathematical relationship $\mathbb{P}(\tilde{y}|\bm{x},\alpha)\mathbb{P}(\tilde{y})=\mathbb{P}(\tilde{y}|\bm{x})\mathbb{P}(\tilde{y}|\alpha)$ for any test data. Since $f_{\bm{\theta}}$ learns this property from the training set, it is crucial that the training set adheres to this equation. To increase the likelihood of this equation and ensure consistency between training on $\mathcal{D}_{p}$ and evaluation during the attack phase, we consider whether modifying $\bm{x}$ to $\bm{x^{\prime}}$ is necessary. This leads to the development of the following theorem.
@@
0:  Benign dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{N}$, surrogate model $f_{\bm{\theta}_{s}}$, poisoning ratio $r$, hyperparameters for UAP-DCT ($\epsilon$, $e_{1}$, $e_{2}$), hyperparameters for TAEs (target $\mu_{y}$, bound $\epsilon_{t}$, step size $\alpha_{t}$, iterations $I_{t}$)
0:  Optimized trigger $\bm{t}$, Modified dataset $\mathcal{D}_{t}$
1:  $\mathcal{D}_{s}\leftarrow$ Randomly sample a subset from $\mathcal{D}$ with ratio $r$
2:  # Train surrogate model and search trigger
3:  $f_{\bm{\theta}_{s}},\bm{t}\leftarrow$ Apply Algorithm 1 with $(\mathcal{D}_{s},f_{\bm{\theta}_{s}},\epsilon,e_{1},e_{2})$
4:  # Generate poisoned subset $\mathcal{D}_{p}$
5:  for each $(\bm{x}_{i},y_{i})\in\mathcal{D}_{s}$ do
6:     $\alpha_{i}\leftarrow(y_{i}-\mu_{y})/\Delta{y}_{t}$
7:     $\bm{x}_{i}^{\prime}\leftarrow$ Targeted-PGD(model=$f_{\bm{\theta}_{s}}$, input=$\bm{x}_{i}$,           target=$\mu_{y}$, bound=$\epsilon_{t}$, step=$\alpha_{t}$, iter=$I_{t}$)
8:  end for
9:  $\mathcal{D}_{p}\leftarrow\{(T(\bm{x}_{i}^{\prime},g(\alpha_{i})\cdot\bm{t}),y_{i})\}_{i=1}^{N_{s}}$
10:  $\mathcal{D}_{c}\leftarrow\mathcal{D}\setminus\mathcal{D}_{s}$, $\mathcal{D}_{t}\leftarrow\mathcal{D}_{c}\cup\mathcal{D}_{p}$
11:  return  $\bm{t}$, $\mathcal{D}_{t}$
@@
Table 2: Imperceptibility of the poisoned data ($\text{PSNR}_{1}\uparrow$).
@@
Datasets and Models. We choose the LIVEC dataset (Ghadiyaram and Bovik 2015) consisting of 1162 images with sizes $500\times 500$ and KonIQ-10k dataset (Hosu et al. 2020) consisting of 10073 images with sizes $512\times 384$. For LIVEC, consistent with prior research (Liu et al. 2024a), we randomly select 80% of the images as the training set, and the remaining 20% as the testing set. For KonIQ-10k, we adopt the official split with 70% of the images as the training set. We include two CNN-based NR-IQA models: HyperIQA (Su et al. 2020), and DBCNN (Zhang et al. 2020), and one Transformer-based TReS (Golestaneh, Dadsetan, and Kitani 2022).
Baseline Attack Methods. We compare our attacks with existing poisoning-based BA. We utilize Blended (Chen et al. 2017), WaNet (Nguyen and Tran 2021), and FTrojan (Wang et al. 2022), which are representative non-patch-based invisible attacks. For all baselines, we employ the coefficient $\alpha$ to adjust the intensity and strength of the triggers, following a similar approach to our attacks. For FTrojan, we add the triggers into the same mid-frequency channels. We apply the same configurations to all methods, except the trigger.
Evaluation Metrics. For the evaluation on benign data, we follow previous works and consider three metrics, i.e., RMSE, PLCC, and SROCC. To evaluate the attack effectiveness, we include the mean absolute error (MAE), and the mean ratio of amplification (MRA) across the test set regarding each $\alpha\!\!\in\!\![-1,1]$. While it is impossible to consider all values of $\alpha$, we estimate it using finite $\alpha$ from a selected set $A$. Therefore, we utilize the mean of those metrics for an overall evaluation:
$\begin{split}&\text{MAE}(\alpha)=\mathop{\mathbb{missing}}{E}\limits_{(\bm{x},y)\sim\mathcal{C}}\Big{[}\big{\lvert}\Delta_{{a}}(\bm{x},\alpha)-\Delta_{{t}}(\alpha)\big{\lvert}\Big{]},\\ &\text{MRA}(\alpha)=\mathop{\mathbb{missing}}{E}\limits_{(\bm{x},y)\sim\mathcal{C}}\Big{[}\frac{\Delta_{{a}}(\bm{x},\alpha)}{\Delta_{{t}}(\alpha)}\Big{]},\\ &\text{mMAE}=\frac{1}{n(A)}\sum_{\alpha\in A}\text{MAE}(\alpha),\text{mMRA}=\frac{1}{n(A)}\sum_{\alpha\in A}\text{MRA}(\alpha),\\ &\text{with}~{}\Delta_{{a}}(\bm{x},\alpha)=f_{\bm{\theta}^{*}}(T(\bm{x},{\alpha}\cdot\bm{t}))-f_{\bm{\theta}^{*}}(\bm{x}),\Delta_{{t}}(\alpha)=\alpha\cdot{\Delta{y}}_{t}.\end{split}$
@@
Settings. As the MOS range is $[0,100]$, we set the maximum deviation $\Delta{y_{t}}=40$, ensuring this deviation is sufficient to yield potent attacks. To regulate the imperceptibility of poisoned images, in Algorithm 1, we set $\lambda=10^{8}$, $e_{1}=24$, $e_{2}=50$, and $\epsilon=\frac{8}{255}$ for LIVEC. For KonIQ-10k, where a larger number of images are eligible for poisoning, we adjust $\epsilon$ to a lower value $\frac{4}{255}$. We set the poisoning ratio $r=20\%$ for all attacks and datasets, and use the HyperIQA as the surrogate model. For the sampling strategy of $\alpha$ to construct $\mathcal{D}_{p}$, we consider sampling from {±$\{\pm 1,\pm\frac{3}{4},\pm\frac{1}{2},\pm\frac{1}{4}\}$ with probabilities $\{0.4,0.3,0.2,0.1\}$, and provide results of other strategies in the ablation study.